OpenStudy (anonymous):

Factor (3x+2y)^3+(y-3)^3

OpenStudy (dumbcow):

sum of cubes: a^3 +b^3 = (a+b)(a^2-ab+b^2)

OpenStudy (anonymous):

I know the formula but i end up getting (3x+3y-3)(9x+4y-3xy-9x+2y^2-6y+y^2-9), and then i simply the first parenthesis which gives me 3(x+y-1) but according to the book it should be 9(x+y-1), and i have no idea how they got that, and i know the second part is wrong too

OpenStudy (dumbcow):

the 9 comes from factoring out another 3 from 2nd part double check the 2nd part (a^2 -ab +b^2) (3x+2y)^2 = 9x^2 +12xy +4y^2 (3x+2y)(y-3) = 3xy-9x-6y+2y^2 (y-3)^2 = y^2 -6y +9

OpenStudy (anonymous):

Well i'm not sure still where the 9 comes from can you show me that........and when i combined like terms for the second part i got (9x^2+9xy+7y^2-12y-9x+9)............is that right, and now i see that i can factor out the 9 in those so that will give me (x^2+xy+7y^2-12y-x+1)........but that's not the answer in the book tho, so i'm not sure where i messed up